OpenStudy (anonymous):

can some one please help me out with an example of the operations on functions, I need an example of an addition problem.....

OpenStudy (anonymous):

\[f(x)=x^2+x\] \[g(x)=3x+1\] \[(f+g)(x)=x^2+x+3x+1=x^2+4x+1\]